# HUNTER: Unsupervised Human-centric 3D Detection via Transferring Knowledge from Synthetic Instances to Real Scenes

### 05 Mar 2024 v1

@@ Abstract
Human-centric 3D scene understanding has recently drawn increasing attention, driven by its critical impact on robotics. However, human-centric real-life scenarios are extremely diverse and complicated, and humans have intricate motions and interactions. With limited labeled data, supervised methods are difficult to generalize to general scenarios, hindering real-life applications. Mimicking human intelligence, we propose an unsupervised 3D detection method for human-centric scenarios by transferring the knowledge from synthetic human instances to real scenes. To bridge the gap between the distinct data representations and feature distributions of synthetic models and real point clouds, we introduce novel modules for effective instance-to-scene representation transfer and synthetic-to-real feature alignment. Remarkably, our method exhibits superior performance compared to current state-of-the-art techniques, achieving a substantial 87.8% improvement in mAP and closely approaching the performance of fully supervised methods (62.15 mAP vs. 69.02 mAP) on HuCenLife.
@@ 1 Introduction
Indeed, humans exhibit an impressive ability to detect objects in 3D spaces based solely on their knowledge of object shapes and sizes, as Fig. 1 shows. Since we can generate arbitrary human instances by the parametric model [20], this prompts us to ponder the following question: “Can our AI method detect human instances in 3D scenes without any scene-level annotations, merely relying on synthetic human models?” To achieve this, we must address three pivotal challenges: the unification of disparate data representations of mesh models and LiDAR point clouds; the alignment of dissimilar feature distributions of synthetic humans and real humans; and the exploitation of prior knowledge specific to the human body to enhance perceptual acuity.
@@ 3.1 Instance-to-Scene Representation Transfer
Range image bridged point cloud generation. However, the simulated LiDAR point cloud only accounts for self-occlusion. To generate scenes that adhere to the view-dependent property of LiDAR point cloud, we simulate the external occlusions by other instances or objects in the scene for inserted humans. Specifically, we utilize the range image $I^{H*W*(N+1)}$ to achieve this, where $H$ and $W$ represents the vertical and horizontal resolution of LiDAR, and $N$ represent the dimension of point cloud (e.g. x, y, z). Due to LiDAR’s limited measuring range, the image has empty grids. Thus, one more dimension is added to represent whether points are present in the grid. If multiple points in the real scene fall into the same grid, LiDAR only records the closest point due to the physical nature of light, which propagates in a straight line. This property forms the external occlusions in real LiDAR-captured data. To simulate the occlusion naturally, we first transform the point cloud of 3D scene and the point cloud of inserted human into the range image $I^{Q}$ and $I^{q}$, respectively. Then, we merge two range images based on the distance $D$ to LiDAR to insert a synthetic human instance into the scene. the formula is presented as
@@
Mask-constrained receptive field control. By using the simulated scene with “annotated” human instances, we can train our model directly. However, as we only “annotated” synthetic humans, those real humans without annotations will be considered as irrelevant backgrounds, inevitably leading the model to ignore real human instances. To address this, we generate a mask to constrain the receptive field to areas where synthetic humans locate while real humans do not.
We choose the renowned anchor-free 3D detection network CenterPoint [44] as our backbone, which gives BEV heatmap and bounding box as outputs, indicating the human’s location and rotation. To generate mask $M$ for vacant ground without objects first, we voxelize the scene and calculate the number of empty voxels above each BEV grid. If the number exceeds a threshold, we regard it as the vacant ground.
During the training phase, this mask $M$ undergoes logical $or$ operation with the ground truth heatmap $y$ created by the synthetic label to generate $M^{*}$, which represents the ground without real humans. Centerpoint [44] utilize Gaussian focal loss for heatmap loss and $\ell_{2}$-norm for hounding box loss. We modify the heatmap loss $L_{hm}$ to work with our receptive control mask:
@@
Note that the mask is only used during training procedure. With “annotated” data and mask-constrained receptive field control, we trained our detector using only the generated synthetic data. We then infer purely on training split to generate pseudo-labels of real humans for the next stage.
@@ 3.2 Synthetic-to-Real Feature Alignment
$L_{s_{2}r}=\left(\overline{F_{s}}-\overline{F_{r}}\right)^{2}$ (4)
@@
$+\frac{1}{\left|F_{r}\right|}\sum_{f_{r}\in F_{r}}R\left(\left|1-\left\|f_{r}\right\|_{2}\right|-\Delta var\right)^{2}$
@@
where $R(\cdot)$ denotes Rectified Linear Unit (ReLU) and $\left\|\cdot\right\|_{2}$ denotes $\ell_{2}$-norm. $\Delta var$ is introduced to regulate the maximum allowable distance. $\beta_{3}$, $\beta_{4}$ are constant weights for loss. With cleaned pseudo-label provided by our bi-directional multi-object filter and BEV feature alignment guiding the model, we retrain the model on our “annotated” scene. After that, the model is inferred on training split again to generate higher-quality pseudo-labels.
@@ 4.1 Experimental setup
Implementation details. We generate synthetic human data using SMPL [19] model with pose and shape parameter from SURREAL [34]. Specifically, we first randomly choose a sequence, then randomly choose a frame within this sequence. After that we put random number of synthetic humans on ground as 3.1 described. For HuCenLife we set the detection range to [25.6, 51.2] meters for the X and Y axis and [-2.5, 7.5] meters for the Z axis. For STCrowd, we focus on the cropped point cloud range with [30.72, 40.96] meters for the X and Y axis and [-4, 1] for the Z axis. Moreover, to fulfill our fine-grained detection necessity, we use denser voxel than the offical setting of STCrowd. Specifically, we use [0.025, 0.05, 0.25] for HuCenLife and [0.03, 0.04, 0.125] for STCrowd. For post-processing of detection results, we use a circle NMS method with a threshold of 0.2 for all the experiments. See more details in the supplementary.
Metrics. Following HuCenLife [41] and STCrowd [10] official metrics, we use Precision, Recall and Average Precision (AP) with 3D center distance threshold $D=\{0.25,0.5,1\}$. Furthermore, we compute the mean Average Precision(mAP), mean Precision (mPrec), mean Recall (mRecall) by averaging AP, Precison and Recall, respectively.
@@ 4.2 Results
Qualitative results.
We provide qualitative results on both HuCenLife and STCrowd in Fig. 3. For OYSTER and DBSCAN, they suffer a lot from false positive detection (shown as orange boxes). Unlike our approach, they are class-agnostic, which consider all the objects after clustering as human, such as walls, television, and even ground that failed to be removed from ground removal process. For MODEST, it relys on motion detector to produce and filter pseudo-labels. In human-centric detection, the challenge arises from the relatively small amplitude of human movements and the possibility that only a portion of the person’s body is in motion. As a result, motion detector provides bounding box with error in size and position in these scenario (shown as black boxes). Furthermore, OYSTER, MODEST, and DBSCAN all suffers from false-negative detection, for the same reason (shown as purple boxes). Despite our method generating false positive detection on the second row, it still outperforms other methods. This truly demonstrates the effectiveness of our approach.
@@ 4.3 Ablation study
Effect of different amount of synthetic data. In this section, we explore the impact of varying amounts of synthetic data on our algorithm. We report the results in Tab. 4. The table suggests that as the number of synthetic frames increases, there is an improvement in performance across the evaluated metrics, demonstrating the positive impact of varying amounts of synthetic data on the algorithm’s effectiveness. In detail, even if we use just 5k synthetic frames (about 27% more than the original annotated training data) for training, the performance still surpass current SOTA methods over 60%. This further demonstrate that our method has a better capacity for capturing human’s both surface level and fine-grained representation. Furthermore, we observed enhancement over all metrics from 5k to 10k, whereas the incremental gains become marginal when we use around 50k synthetic frames. This observation demonstrates that increasing the number of synthetic frames improve performance but the payback start to diminish as frame number increase. Hence, we selected 50k frames to strike the balance between performance and efficiency for training.
@@ 4.4 Effect of feature extractor
We evaluate our feature extractor’s effectiveness in Tab. 5 by fine-tuning with varying proportions of ground truth data. Note that we freeze all the parameters and weights for the feature extractor.
Remarkably, our feature extractor’s performance exceeds that of the fully supervised approach by a substantial margin, when utilizing an equivalent amount of ground truth data. With only 10% of ground truth data employed for fine-tuning, our feature extractor surpasses the performance of the fully supervised method trained on 100% ground truth data. Furthermore, there is a notable 4.7% improvement over the fully supervised method when utilizing the full 100% ground truth data for fine-tuning. This noteworthy result further underscores the efficacy of our feature extractor, showcasing its ability to adapt and refine feature representations.
@@ 5 Conclusions
We introduce an unsupervised 3D detection approach for human-centric scenarios, which transfers knowledge from synthetic human models to actual scenes. We develope effective modules for instance-to-scene representation transfer and synthetic-to-real feature alignment to overcome the disparities in data representations and feature distributions between synthetic models and real-world point clouds. Impressively, our method outperforms existing state-of-the-art techniques, achieving a significant 87.8% improvement in mAP, and nearly matching the results of fully supervised methods. Looking ahead, we plan to explore the method’s applicability with temporal features and other scenarios.
@@ Appendix C More result visualization
Fine-Grained Perception Enhancement is designed to ease the occlusion problem. $E_{7}$ in Tab. 3 has shown its effectiveness for the whole data with self- or external occlusions. We show the improvement from $E_{6}$ to $E_{7}$ for occluded instances with Fig. 6 on HCL test set.

### 15 Mar 2024 v2

@@ Abstract
Human-centric 3D scene understanding has recently drawn increasing attention, driven by its critical impact on robotics. However, human-centric real-life scenarios are extremely diverse and complicated, and humans have intricate motions and interactions. With limited labeled data, supervised methods are difficult to generalize to general scenarios, hindering real-life applications. Mimicking human intelligence, we propose an unsupervised 3D detection method for human-centric scenarios by transferring the knowledge from synthetic human instances to real scenes. To bridge the gap between the distinct data representations and feature distributions of synthetic models and real point clouds, we introduce novel modules for effective instance-to-scene representation transfer and synthetic-to-real feature alignment. Remarkably, our method exhibits superior performance compared to current state-of-the-art techniques, achieving 87.8% improvement in mAP and closely approaching the performance of fully supervised methods (62.15 mAP vs. 69.02 mAP) on HuCenLife Dataset.
@@ 1 Introduction
Indeed, humans exhibit an impressive ability to detect objects in 3D spaces based solely on their knowledge of object shapes and sizes, as Fig. 1 shows. Since we can generate arbitrary human instances by the parametric model [20], this prompts us to ponder the following question: “Can our AI method detect human instances in 3D scenes without any scene-level annotations, merely relying on synthetic human models?" To achieve this, we must address three pivotal challenges: the unification of disparate data representations of mesh models and LiDAR point clouds; the alignment of dissimilar feature distributions of synthetic humans and real humans; and the exploitation of prior knowledge specific to the human body to enhance perceptual acuity.
@@ 3.1 Instance-to-Scene Representation Transfer
Range image bridged point cloud generation. However, the simulated LiDAR point cloud only accounts for self-occlusion. To generate scenes that adhere to the view-dependent property of LiDAR point cloud, we simulate the external occlusions by other instances or objects in the scene for inserted humans. Specifically, we utilize the range image $I^{H*W*(N+1)}$ to achieve this, where $H$ and $W$ represents the vertical and horizontal resolution of LiDAR, and $N$ represent the dimension of point cloud (e.g. x, y, z). Due to LiDAR’s limited measuring range, the image has empty grids. Thus, one more dimension is added to represent whether points are present in the grid. If multiple points in the real scene fall into the same grid, LiDAR only records the closest point due to the physical nature of light, which propagates in a straight line. This property forms the external occlusions in real LiDAR-captured data. To simulate the occlusion naturally, we first transform the point cloud of 3D scene and the point cloud of inserted human into the range image $I^{Q}$ and $I^{q}$, respectively. Then, we merge two range images based on the distance $D$ to LiDAR to insert a synthetic human instance into the scene. the formula is
@@
Mask-constrained receptive field control. By using the simulated scene with “annotated" human instances, we can train our model directly. However, as we only “annotated" synthetic humans, those real humans without annotations will be considered as irrelevant backgrounds, inevitably leading the model to ignore real human instances. To address this, we generate a mask to constrain the receptive field to areas where synthetic humans locate while real humans do not. We choose the renowned anchor-free 3D detection network CenterPoint [44] as our backbone, which gives BEV heatmap and bounding box as outputs, indicating the human’s location and rotation. To generate mask $M$ for vacant ground without objects first, we voxelize the scene and calculate the number of empty voxels above each BEV grid. If the number exceeds a threshold, we regard it as the vacant ground. During the training phase, this mask $M$ undergoes logical $or$ operation with the ground truth heatmap $y$ created by the synthetic label to generate $M^{*}$, which represents the ground without real humans. Centerpoint [44] utilize Gaussian focal loss for heatmap loss and $\ell_{2}$-norm for hounding box loss. We modify the heatmap loss $L_{hm}$ to work with our receptive control mask:
@@
Note that the mask is only used during training procedure. With “annotated" data and mask-constrained receptive field control, we trained our detector using only the generated synthetic data. We then infer purely on training split to generate pseudo-labels of real humans for the next stage.
@@ 3.2 Synthetic-to-Real Feature Alignment
$L_{s_{2}r}=\left(\overline{F_{s}}-\overline{F_{r}}\right)^{2},$ (4)
@@
$+\frac{1}{\left|F_{r}\right|}\sum_{f_{r}\in F_{r}}R\left(\left|1-\left\|f_{r}\right\|_{2}\right|-\Delta var\right)^{2},$
@@
where $R(\cdot)$ denotes Rectified Linear Unit (ReLU) and $\left\|\cdot\right\|_{2}$ denotes $\ell_{2}$-norm. $\Delta var$ is introduced to regulate the maximum allowable distance. $\beta_{3}$, $\beta_{4}$ are constant weights for loss. With cleaned pseudo-label provided by our bi-directional multi-object filter and BEV feature alignment guiding the model, we retrain the model on our “annotated" scene. After that, the model is inferred on training split again to generate higher-quality pseudo-labels.
@@ 4.1 Experimental setup
Implementation details. We generate synthetic human data using SMPL [19] model with pose and shape parameter from SURREAL [34]. Specifically, we first randomly choose a sequence, then randomly choose a frame within this sequence. After that we put random number of synthetic humans on ground as 3.1 described. For HuCenLife we set the detection range to [25.6, 51.2] meters for the X and Y axis and [-2.5, 7.5] meters for the Z axis. For STCrowd, we focus on the cropped point cloud range with [30.72, 40.96] meters for the X and Y axis and [-4, 1] for the Z axis. Moreover, to fulfill our fine-grained detection necessity, we use denser voxel than the offical setting of STCrowd. Specifically, we use [0.025, 0.05, 0.25] for HuCenLife and [0.03, 0.04, 0.125] for STCrowd. For post-processing of detection results, we use a circle NMS method with a threshold of 0.2 for all the experiments.
Metrics. Following HuCenLife [41] and STCrowd [10] official metrics, we use Precision, Recall and Average Precision (AP) with 3D center distance threshold $D=\{0.25,0.5,1\}$. Furthermore, we compute the mean Average Precision(mAP), mean Precision (mPrec), mean Recall (mRecall) by averaging AP, Precison and Recall.
@@ 4.2 Results
Qualitative results. We provide qualitative results on both HuCenLife and STCrowd in Fig. 3. For OYSTER and DBSCAN, they suffer a lot from false positive detection (shown as orange boxes). Unlike our approach, they are class-agnostic, which consider all the objects after clustering as human, such as walls, television, and even ground that failed to be removed from ground removal process. For MODEST, it relys on motion detector to produce and filter pseudo-labels. In human-centric detection, the challenge arises from the relatively small amplitude of human movements and the possibility that only a portion of the person’s body is in motion. As a result, motion detector provides bounding box with error in size and position in these scenario (shown as black boxes). Furthermore, OYSTER, MODEST, and DBSCAN all suffers from false-negative detection, for the same reason (shown as purple boxes). Despite our method generating false positive detection on the second row, it still outperforms other methods. This truly demonstrates the effectiveness of our approach.
@@ 4.3 Ablation study
Effect of different amount of synthetic data. In this section, we explore the impact of varying amounts of synthetic data on our algorithm. We report the results in Tab. 4. The table suggests that as the number of synthetic frames increases, there is an improvement in performance across the evaluated metrics, showing the positive impact of varying amounts of synthetic data on the algorithm’s effectiveness. In detail, even if we use just 5k synthetic frames (about 27% more than the original annotated training data) for training, the performance still surpass current SOTA methods over 60%. This further demonstrate that our method has a better capacity for capturing human’s both surface level and fine-grained representation. Furthermore, we observed enhancement over all metrics from 5k to 10k, whereas the incremental gains become marginal when we use around 50k synthetic frames. This observation demonstrates that increasing the number of synthetic frames improve performance but the payback start to diminish as frame number increase. Hence, we selected 50k frames to strike the balance between performance and efficiency for training.
@@ 4.4 Effect of feature extractor
We evaluate our feature extractor’s effectiveness in Tab. 5 by fine-tuning with varying proportions of ground truth data. Note that we freeze all the parameters and weights for the feature extractor. Remarkably, our feature extractor’s performance exceeds that of the fully supervised approach by a substantial margin, when utilizing an equivalent amount of ground truth data. With only 10% of ground truth data employed for fine-tuning, our feature extractor surpasses the performance of the fully supervised method trained on 100% ground truth data. Furthermore, there is a notable 4.7% improvement over the fully supervised method when utilizing the full 100% ground truth data for fine-tuning. This noteworthy result further underscores the efficacy of our feature extractor, showcasing its ability to adapt and refine feature representations.
@@ 5 Conclusions
We introduce an unsupervised 3D detection approach for human-centric scenarios, which transfers knowledge from synthetic human models to actual scenes. We develop effective modules for instance-to-scene representation transfer and synthetic-to-real feature alignment to overcome the disparities in data representations and feature distributions between synthetic models and real-world point clouds. Impressively, our method outperforms existing state-of-the-art methods with a significant margin, and nearly matching the results of fully supervised methods.
@@ Appendix C More result visualization
Fine-Grained Perception Enhancement is designed to ease the occlusion problem. $E_{7}$ in ablation study has shown its effectiveness for the whole data with self- or external occlusions. We show the improvement from $E_{6}$ to $E_{7}$ for occluded instances with Fig. 6 on HuCenLife test set.